OpenStudy (anonymous):

what are the whole number equivalents of 10/5 besides 2?

jimthompson5910 (jim_thompson5910):

10/5 = 2, and this is the only whole number that it's equal to

OpenStudy (anonymous):

ok but the worksheet says equivalents ok thx

jimthompson5910 (jim_thompson5910):

well 10/5 = 2/1 and 10/5 = 4/2

jimthompson5910 (jim_thompson5910):

also, 10/5=6/3 and 10/5=14/7

OpenStudy (anonymous):

ok how about 20/4 equivalents besides 5 ?

jimthompson5910 (jim_thompson5910):

20/4 = 10/2 20/4 = 15/3 20/4 = 25/5 20/4 = 30/6